OpenStudy (anonymous):

Finish writing the equation of the line y = -1/2x + ?, given the points (8, -3) and (-2, 2).

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

find the ? for line in which it intersects both points

OpenStudy (anonymous):

y=-1/2x ?

OpenStudy (anonymous):

y = -1/2x + ? 1/2x+ y = -1/2x + ? + 1/2x 1/2x+ y = ? use point (-2,2) x=-2 and y=2 1/2*(-2)+ 2 = ? -1 + 2 = ? = 1